# Exercising the CCPA Opt-out Right on Android: Legally Mandated but Practically Challenging

### v4

@@ 7.2. Two Settings for One Task
As it stands, Android currently has two types of opt-out settings: the platform-level AdID and the app-level CCPA opt-out settings. According to Google, if a user has disabled the AdID, operators “may not use the advertising identifier for creating user profiles for advertising purposes or for targeting users with personalized advertising” (Google, 2024c). Similarly, the CCPA prohibits the sharing of personal information for purposes of cross-context behavioral advertising, defined as the “targeting of advertising to a consumer based on the consumer’s personal information” and which covers “consumer profiles” per the CCPA Regulations. [^26] CCPA §1798.140(k), CCPA Regulations, §7025(c)(1). Thus, the intended effect of the AdID setting is similar to what the CCPA opt-out right is intended to accomplish. However, despite this similarity the AdID setting is without legal meaning outside the contractual relationship that Google has with the app operators.
@@ 7.3. Re-purposing the AdID Setting
Disabling the AdID should be interpreted as exercising the opt-out right. This would immediately make the opt-out right on Android available to a wide audience in line with previous results suggesting that many users have interest in such setting (Zimmeck et al., 2023). Such re-purposing would unify and streamline the current two types of settings—the platform-level AdID setting and the app-level opt-out right settings—into one and lead to greater legal clarity. The current AdID API can be modified to return an opt-out signal (akin to GPC’s Sec-GPC: 1 header) instead of a string of zeros when a user opts out (Google, 2024d). Notably, opting out per the CCPA via GPC and enabling the AdID setting are both intended as universal opt-out settings and have nearly identical scope (Appendix 10.2). However, ultimately, whether this uniform setting is a GPC setting or other opt-out right setting is secondary.
@@
Exercising the opt-out right may have impact on other platform-level privacy settings, which would need to be aligned. Notably, while its further direction is not yet fully clear (Chavez, 2025), Google’s Privacy Sandbox, a privacy framework for the web and Android, includes various settings that could be impacted. In general, if users exercise their opt-out right under the CCPA, any setting that would enable the selling or sharing of personal information would need to be turned off. The opt-out right setting should also be available to users on devices without Google Play Services, i.e., on alternative Android distributions. As the simplest solution, those could re-implement the Google APIs as, e.g., GrapheneOS (GrapheneOS, 2024) and /e/OS (e Foundation, 2024), already do. Overall, the emerging legal landscape requiring online ads to be served in a privacy-preserving manner presents an opportunity for Google to re-purpose the AdID setting towards the opt-out right and align them with the mandatory legal requirements in California and other jurisdictions.
@@ 8. Conclusions
We study the effectiveness of opting out per the CCPA from ad tracking via the Android AdID setting, GPC, and app-level opt-out UIs. We find that the CCPA opt-out right applies to 62%–81% of the top-free Android apps on the US Play Store and that many apps are not compliant. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out. The main problem is that the opt-out right is not supported by a platform-level setting. Moving forward, as the AdID setting is not aligned with users’ opt-out right per the CCPA, the current AdID setting should be re-purposed towards an opt-out right setting. As we focused on Android, the opt-out compliance on iOS and other platforms remains an open question. It may be that the App Tracking Transparency framework (Apple Developer, 2024) does not help users with their opt-out right either resulting in similar compliance rates as we found here for Android. In general, it would be a worthwhile endeavor to evaluate the extent to which opt-out rights are honored on different platforms and for different laws.
@@ 9. Ethical Principles
We designed and conducted our study to minimize risk and adhere to ethical research principles. All data collection was limited to the personal information of the researchers. As the study did not involve external human subjects, it did not require review by our institutions’ Institutional Review Board. The impact on the evaluated apps’ infrastructure was minimal as it was limited to approximately 30 seconds per app. In line with responsible disclosure principles, we have notified all companies mentioned by name in this paper whose services were observed accessing the Android AdID after it had been disabled. We are also in the process of contacting app operators who stated that the CCPA opt-out right is available to their users in their privacy policy but failed to provide a corresponding setting. The use of Mitmproxy for the purpose of our research is a standard and common methodology (Kollnig et al., 2022a).

### v5

@@ 7.2. Two Settings for One Task
As it stands, Android currently has two types of opt-out settings: the platform-level AdID and the app-level CCPA opt-out settings. The AdID setting is not intended for exercising the opt-out right per the CCPA or other privacy laws. However, according to Google, if a user has disabled the AdID, operators “may not use the advertising identifier for creating user profiles for advertising purposes or for targeting users with personalized advertising” (Google, 2024c). Similarly, the CCPA prohibits the sharing of personal information for purposes of cross-context behavioral advertising, defined as the “targeting of advertising to a consumer based on the consumer’s personal information” and which covers “consumer profiles” per the CCPA Regulations. [^28] CCPA §1798.140(k), CCPA Regulations, §7025(c)(1). Thus, the intended effect of the AdID setting is similar to what the CCPA opt-out right is intended to accomplish. However, despite this similarity the AdID setting is without legal meaning outside the contractual relationship that Google has with the app operators.
@@ 7.3. Re-purposing the AdID Setting
Disabling the AdID should be interpreted as exercising the opt-out right. This would immediately make the opt-out right on Android available to a wide audience in line with previous results suggesting that many users have interest in such setting (Zimmeck et al., 2023). Such re-purposing would unify and streamline the current two types of settings—the platform-level AdID setting and the app-level opt-out right settings—into one and lead to greater legal clarity. The current AdID API can be modified to return an opt-out signal (akin to GPC’s Sec-GPC: 1 header) instead of a string of zeros when a user opts out (Google, 2024d). Notably, opting out via GPC and the AdID setting are both intended as universal settings and have nearly identical scope (Appendix 10.2). However, ultimately, whether this uniform setting is a GPC setting or other opt-out right setting is secondary.
@@
Exercising the opt-out right may have impact on other platform-level privacy settings, which would need to be aligned. Notably, while its further direction is not yet fully clear (Chavez, 2025), Google’s Privacy Sandbox, a privacy framework for the web and Android, includes various settings that could be impacted. In general, if users exercise their opt-out right, any setting that would run contrary to what the right requires would need to be turned off, e.g., under the CCPA, settings that would enable the selling or sharing of personal information. The opt-out right setting should also be available to users on devices without Google Play Services, i.e., on alternative Android distributions. As the simplest solution, those could re-implement the Google APIs as, e.g., GrapheneOS (GrapheneOS, 2024) and /e/OS (e Foundation, 2024), already do. Overall, the emerging legal landscape requiring online ads to be served in a privacy-preserving manner presents an opportunity for Google to re-purpose the AdID setting towards the opt-out right and align them with the mandatory legal requirements in California and other jurisdictions.
@@ 8. Conclusions
We study the effectiveness of opting out per the CCPA from ad tracking via GPC and app-level opt-out UIs. We find that the CCPA opt-out right applies to 62%–81% of the top-free Android apps on the US Play Store and that many apps are not compliant. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out. The main problem is that the opt-out right is not supported by a platform-level setting. Moving forward, as the AdID setting is not aligned with users’ opt-out right, the current AdID setting should be re-purposed towards an opt-out right setting. As we focused on Android, the opt-out compliance on iOS and other platforms remains an open question. It may be that the App Tracking Transparency framework (Apple Developer, 2024) does not help users with their opt-out right either resulting in similar compliance rates as we found here for Android. In general, it would be a worthwhile endeavor to evaluate the extent to which opt-out rights are respected on different platforms and for different laws.
@@ 9. Ethical Principles
We designed and conducted our study to minimize risk and adhere to ethical research principles. All data collection was limited to the personal information of the researchers. As the study did not involve external human subjects, it did not require review by our institutions’ Institutional Review Board. The impact on the evaluated apps’ infrastructure was minimal as it was limited to approximately 30 seconds per app. In line with responsible disclosure principles, we have notified all companies mentioned by name in this paper whose services were observed accessing the Android AdID after it had been disabled. We are also in the process of contacting app operators who stated that the CCPA opt-out right is available to California residents in their privacy policy but failed to provide a corresponding setting. The use of Mitmproxy for the purpose of our research is a standard and common methodology (Kollnig et al., 2022a).